OpenStudy (anonymous):

Which of these has the greatest area? Use three point one four for pi. rhombus with side lengths: 5 inches, height: 6 inches triangle with base length: ten inches, height: sixteen inches circle with diameter: 5 inches trapezoid with base lengths: 9 inches and eleven inches, height: 6 inches

OpenStudy (anonymous):

rhombus area = 5*6 = 30 triangle area = 1/2*10*16 = 80 circle area = pi*r^2 = pi*(d/2)^2 = pi*(2.5)^2 = approx 20 trapezoid area, if I recall correctly is: \[\frac{1}{2}(base1+base2)*(height)\] = 60 The units of all the answers are in square inches